Sodium chloride is produced from its elements through synthetic reaction. What mass of each reactant would be required to produce 25.0 mol of sodium chloride



1
Expert's answer
2022-05-09T17:01:06-0400

Solution:

Balanced chemical equation:

2Na + Cl2 → 2NaCl

According to stoichiometry:

2 mol of Na react with 1 mol of Cl2 to produce 2 mol of NaCl

Thus,

Moles of Na = (25.0 mol NaCl) × (2 mol Na / 2 mol NaCl) = 25.0 mol Na

Moles of Cl2 = (25.0 mol NaCl) × (1 mol Cl2 / 2 mol NaCl) = 12.5 mol Cl2


The molar mass of Na is 22.99 g/mol

Therefore,

Mass of Na = (25.0 mol Na) × (22.99 g Na / 1 mol Na) = 574.75 g Na

Mass of Na = 574.75 g

The molar mass of Cl2 is 70.906 g/mol

Therefore,

Mass of Cl2 = (12.5 mol Cl2) × (70.906 g Cl2 / 1 mol Cl2) = 886.325 g Cl2

Mass of Cl2 = 886.325 g


Answer:

The mass of Na is 574.75 grams

The mass of Cl2 is 886.325 grams